a boy starts walking to school @ 7:30 am at the rate of 3miles per hour, at 7:40am his brother starts for school on his bicycle at 6 miles per hour. When will the boy on the bicycle overtake his brother?
Answer by lwsshak3(11628)

a boy starts walking to school @ 7:30 am at the rate of 3miles per hour, at 7:40am his brother starts for school on his bicycle at 6 miles per hour. When will the boy on the bicycle overtake his brother
**
let x=hrs it takes for boy bicycling to school
10 min=1/6 hr
x+1/6=hrs it takes for boy walking to school
..
distance=speed*travel time (same for both)
6x=3(x+1/6)
6x=3x+3/6
3x=1/2
x=1/6 hr
x+1/6=2/6=1/3 hr
Boy on the bicycle overtakes his brother walking in 1/6 hr or 10 min at 7:50 am